OpenStudy (anonymous):

HELP! i just need help figuring out how to solve this i think i did it wrong could someone check and let me know?

OpenStudy (anonymous):

Show all work to simplify 2 over x + 2 over x minus 1 - 2 over x minus 2. Use complete sentences to explain how to simplify this expression. Remember to list all restrictions.

OpenStudy (anonymous):

my work: first we need a common so i multiply the first epression by x^2-3x+2, the second expression by x^2-2x and the third expression by x^2-x. then i cancel out all the denominators and now i have 2x^2-6x+4+2x^2-4x-2x^2-2x so now i simplify and add like terms 2x^2-12x+4 then i factor out the gcf 2(x^2-6x+2) and thats out final answer restriction= 0,1,2

OpenStudy (anonymous):

I CAN HELP U

OpenStudy (anonymous):

is that right or wrong? i feel like this is wrong

OpenStudy (ankit042):

hey @swaggie can you write/draw the original equation...it is a bit confusing

OpenStudy (anonymous):

Its wrong

OpenStudy (anonymous):

@ankit042 go away I got it

OpenStudy (anonymous):

@swaggie please write out your equation mam

OpenStudy (ankit042):

@snewton29 sure you help her :D

OpenStudy (anonymous):

@ankit042 Seriously man GO AWAY

OpenStudy (anonymous):

\[\frac{ 2 }{ x }+\frac{ 2 }{ x-1 }-\frac{ 2 }{ x-2 }\]

OpenStudy (anonymous):

better?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what's that?^^

OpenStudy (anonymous):

their common denominator is x(x-1)(x-2) so multiply each on by that

OpenStudy (anonymous):

no one asked u @CrazyAsianTracy

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

@snewton29 please be nice

OpenStudy (anonymous):

o

OpenStudy (anonymous):

@CrazyAsianTracy i already did that

OpenStudy (anonymous):

no

OpenStudy (anonymous):

?

OpenStudy (primeralph):

What?

OpenStudy (ankit042):

what is your problem @snewton29 ? You are not even helping and not letting others to contribute

OpenStudy (primeralph):